ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

11. Mathematical Physics

1991 MS Classification: X Bănulescu, Martha, Bucharest, Romania Lie Groups in Physics In the first part the paper we recall our results of Potential Theory obtained for Lie Groups, in general, and on Nilpotent Lie Groups, in particular. In the second part of our paper, we give examples from Physics: the give concrete realizations of the concepts that have been introduced. The recent developments have taken physicists increasingly deeper into the fascinating world of the pure Mathematics.